OpenStudy (anonymous):

Substitue I=100 r=0.01 t=1/2 into the formula for simple intrest and solve for P I=Prt 100 =P(0.01) (1/2) 100= P(0.05) 2,000=P multiply and solve for p I know how to get everything but the 2,000. Need the forumla on what they are multiply and how they are multiply to get the 2,000

myininaya (myininaya):

let see you divided both sides by 0.05 giving you 100=P*(1/100)(1/2) 100=P*(1/200) 2000=P by multiplying 200 on both sides

OpenStudy (anonymous):

still kinda lost sorry

OpenStudy (anonymous):

or what would problom look like when divied

myininaya (myininaya):

do you know .01=1/100?

myininaya (myininaya):

Examples: 1/10=.1, 1/100=.01, 1/1000, .001, 1/10000=.0001 so on....

OpenStudy (anonymous):

yea i know that

myininaya (myininaya):

ok do you know how to multiply 1/100 * 1/2

myininaya (myininaya):

=(1*1)/(100*2)=1/200

myininaya (myininaya):

to multiply fractions you multiply straight across on the numerator and then for the denominator like I did above

myininaya (myininaya):

Since we are trying to get P by itself and it is being multiplied by that 1/200 to undo this multiplication we multiply by the multiplicative inverse of 1/200 to get the multiplicative identity 1. The multiplicative inverse of 1/200 is just its reciprocal which is 200. So we multiplied by 200 to get the multiplicative identity. The multiplicative identity times any number will be that number since the multiplicative identity is 1. But since we multiply 200 on one side we have to do it to the other side of the equation

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

nope thanks alot

myininaya (myininaya):

really you're good?

myininaya (myininaya):

isn't the word multiplicative cool?

myininaya (myininaya):

say yes. you know you want to

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol, random

myininaya (myininaya):

i had the hardest time saying that word, and I finally got it. I used to couldn't say it

OpenStudy (anonymous):

lol well not really i man i kinda get the problem but i mean

myininaya (myininaya):

I like to change everything to fractions, but you could have left it "deciamly"

myininaya (myininaya):

decimaly*

OpenStudy (anonymous):

yea i show you next problem and then show you what i think then can maybe walk me throw it lol like i have done others on here before lol

OpenStudy (anonymous):

i will show you problem

myininaya (myininaya):

k bring it

OpenStudy (anonymous):

there is a screen shoot of the problem

OpenStudy (anonymous):

i know that the fraction will be 3/4

OpenStudy (anonymous):

because 9/12=0.075 wich is 3/4 in fraction

myininaya (myininaya):

are you talking about your P is 3/4? If so, then thats incorrect

OpenStudy (anonymous):

so i will take 0.04 (3/4)

OpenStudy (anonymous):

t

OpenStudy (anonymous):

becasue be 9 /12

myininaya (myininaya):

Our formula for interest is I=Prt where r=4% and t=9 and I=42 and they ask us to solve for P. Let me see you plug those in

OpenStudy (anonymous):

okay give me sec

myininaya (myininaya):

k

OpenStudy (anonymous):

I=42,r=0.04,and t=9/12=3/4 thats what i got so far

OpenStudy (anonymous):

and in terms i will have

OpenStudy (anonymous):

I=prt 42=(0.04)(3/4) 42=P(0.03) then how will i caculate to get P

myininaya (myininaya):

I thought t=9? let me look again

myininaya (myininaya):

t=9 not 3/4

OpenStudy (anonymous):

i will show you an exapmle it gives me how to do problem

OpenStudy (anonymous):

same problem with exapmle beside it

OpenStudy (anonymous):

but say's got to divide 9 by 12 months

myininaya (myininaya):

oh i didnt see the month thing sorry

myininaya (myininaya):

yes there are 12 months so the time will be 9/12 gj

OpenStudy (anonymous):

wich means will be 0.075 in dec and 3/4 in fraction

myininaya (myininaya):

3/4=.75

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I=prt 42=(0.04)(3/4) 42=P(0.03) do i got to divide the 0.04 and (3/4)

myininaya (myininaya):

yes so far correct

OpenStudy (anonymous):

but if i divide i will get 0.05

OpenStudy (anonymous):

if i divide 0.04 and 3/4

OpenStudy (anonymous):

get 0.05 right

myininaya (myininaya):

since we should be multiplying those we get .03

OpenStudy (anonymous):

all i get it

OpenStudy (anonymous):

now thats so crazy lol

OpenStudy (anonymous):

then i take 0.03/42 and that =P

OpenStudy (anonymous):

i mean *

myininaya (myininaya):

42/.03

OpenStudy (anonymous):

=140

OpenStudy (anonymous):

thats crazy i been working on problem couple days and all i need to do lol

myininaya (myininaya):

1400

OpenStudy (anonymous):

i knew how to work problem

OpenStudy (anonymous):

thanks for all your help man thanks alot

myininaya (myininaya):

ur welcome

OpenStudy (anonymous):

maybe can help me with couple more

OpenStudy (anonymous):

i wirked on i got most of problem but can figure certin ones out

OpenStudy (anonymous):

parts of them

myininaya (myininaya):

k

OpenStudy (anonymous):

like i get 33.25 ______ 1900 (3/13) make r=

OpenStudy (anonymous):

3/12

myininaya (myininaya):

yes looks good

myininaya (myininaya):

the first one does. did you want me to check the other two or something?

myininaya (myininaya):

what was your strategy for the next one?

OpenStudy (anonymous):

the frist two are good what i am having trouble with is the intrest rate

OpenStudy (anonymous):

actual intrest

myininaya (myininaya):

what is the formula for actual interest?

OpenStudy (anonymous):

look at the example shows how to get the actual interrest

OpenStudy (anonymous):

i know how to do problem but like when i divide then top by botton and add the bottem times ther fraction not sure i am doing ti right

myininaya (myininaya):

so r=I/(Pt)=33.25/(1866.75*3/12)

myininaya (myininaya):

i will bbl maybe

OpenStudy (anonymous):

lol